OpenStudy (anonymous):

A candy bar box is in the shape of a triangular prism. The volume of the box is 3,240 cubic centimeters.

OpenStudy (anonymous):

Part A: What is the height of the box? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make this candy box? Explain how you got your answer.

OpenStudy (anonymous):

@dumbcow @Astrophysics

OpenStudy (misssunshinexxoxo):

I'd say follow the formula. @jim_thompson5910

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

help :(!!!!

OpenStudy (mathstudent55):

Do you know how to find the volume of a prism?

OpenStudy (mathstudent55):

\(\large V_{prism} = Bh\) where B = area of the base h = height of the prism

OpenStudy (mathstudent55):

The volume of the prism is the area of the base of the prism, a triangle, multiplied by the height of the prism.

OpenStudy (anonymous):

ok so what would my answer be for both questions? @mathstudent55

OpenStudy (mathstudent55):

I don't know. I only know the answer after I work it out.

OpenStudy (mathstudent55):

We can find the area of the base of the prism. We also know the volume of the prism, so we can find the height of the prism. \(V = Bh\) \(3240~cm^3 = \dfrac{24~cm \times 15~cm}{2} \times h \) Solve for h.